Karen walsh has twice as much money invested at 5% simple interest as she does at 4%. If her yearly income from these two investments is $350, how much does she have invested at each rate.
Found 2 solutions by mananth, ikleyn:
Answer by mananth(16949)

let the amount invested at4% be x
The amount invested at 5%=2x
The interest obtained = 350 for 1/2 year or 700 per year
0.04x+0.05*2x= 700
0.04x+0.1x=700
0.14 x =700
x=700/0.14
x=5000 which is the amount invested at4%
2*5000 = 10000 is the amount invested at5%




Answer by ikleyn(53937)

.
Karen Walsh has twice as much money invested at 5% simple interest as she does at 4%.
If her yearly income from these two investments is $350, how much does she have invested at each rate.
~~~~~~~~~~~~~~~~~~~~~~~~


        In the post by @mananth, the condition is interpreted incorrectly.
        As a consequence to it, the setup equation is not adequate to the problem,
        as well as the solution and the answer.

        I came to bring a correct solution.


let the amount invested at 4% be x
The amount invested at 5%i s 2x dollars
The yearly interest obtained is 350 dollars.
0.04x + 0.05*(2x) = 350
0.04x + 0.1x = 350
0.14x = 350
x = 350/0.14
x = 2500 which is the amount invested at 4%
2*2500 = 5000 dollars is the amount invested at 5%

CHECK.     0.04*2500 + 0.05*5000 = 350 dollars.     ! Precisely correct !

Solved correctly.